OpenStudy (anonymous):

The points (1,-2) and (-4,10) are adjacent vertices of a square. What is the perimeter of the square?

OpenStudy (anonymous):

You can use the same approach that we used before, except you won't be squaring the length that you get, you will be adding it to itself 4 times, or just multiplying it by 4 since the perimeter of a square of side + side + side + side, or 4 * side.

OpenStudy (anonymous):

\[\sqrt{(-4 - 1)^2 + (10 - -2)^2}\]

OpenStudy (anonymous):

\[4 * \sqrt{25 + 144}\]

OpenStudy (anonymous):

4 * 13 = 52 units